OpenStudy (anonymous):

the occupations of biomedical engineers, skin care specialists, and physician assistants are among the 15 with the largest growth from 2008 to 2018. the number of physician assistant jobs will grow 7 thousand less than 3 times the number of biomedical engineer jobs. the number of skin care specialist jobs will grow 12 thousand more than half the number of biomedical engineer jobs. If the total growth of these three jobs is predicted to be 59 thousand, find the predicted growth of each job. The predicted growth of biomedical engineers is ___ thousands The predicted growth of skin care specialists is ____ thousands The predicted growth of physician assistants is ____ thousands

OpenStudy (anonymous):

can anyone help? I've been working on this forever now...

OpenStudy (anonymous):

is there anyone out there?

OpenStudy (anonymous):

I need help anyone...

jimthompson5910 (jim_thompson5910):

Let B = The predicted growth of biomedical engineers S = The predicted growth of skin care specialists P = The predicted growth of physician assistants all of these values will be in thousands (ex: saying B = 2 means the number of biomed engineers grew by 2 thousand)

jimthompson5910 (jim_thompson5910):

"the number of physician assistant jobs will grow 7 thousand less than 3 times the number of biomedical engineer jobs" which means P = 3B - 7

jimthompson5910 (jim_thompson5910):

"the number of skin care specialist jobs will grow 12 thousand more than half the number of biomedical engineer jobs" telling us that S = (1/2)B + 12

jimthompson5910 (jim_thompson5910):

" the total growth of these three jobs is predicted to be 59 thousand" so B+S+P = 59

jimthompson5910 (jim_thompson5910):

making sense so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

we can replace all of S with (1/2)B + 12 we can replace all of P with 3B - 7

jimthompson5910 (jim_thompson5910):

doing those two things makes us go from B+S+P = 59 to B+(1/2)B + 12 + 3B - 7 = 59

jimthompson5910 (jim_thompson5910):

do you see how to solve for B?

OpenStudy (anonymous):

uh im little lost

jimthompson5910 (jim_thompson5910):

where at?

OpenStudy (anonymous):

sorry i was working it out on paper lol, so B=12

OpenStudy (anonymous):

B=12?

OpenStudy (anonymous):

are you still there?

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

yes I'm getting B = 12 as well

jimthompson5910 (jim_thompson5910):

using B = 12, you can figure out the values of S and P using P = 3B - 7 S = (1/2)B + 12

OpenStudy (anonymous):

so replace all "B' with 12

OpenStudy (anonymous):

P=29 and S=18 is that right?

jimthompson5910 (jim_thompson5910):

yes and then evaluate

jimthompson5910 (jim_thompson5910):

correct, I'm getting the same for P and S

OpenStudy (anonymous):

so the biomedical engineer is 12 thousand the skin care specialists is 18 thousand the physician assistants is 29 thousand

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

YAHOO!!!! Thank you so much I have been stuck on this for like 30 minutes now, again thank you :), you have a new fan

jimthompson5910 (jim_thompson5910):

you're welcome